Session: Oct 2005 (session 3)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Marco Poloni at Wed Oct 26 09:30:02 2005) Success :
Experiment with foggy conditions

Urumqi (Zhang hua at Wed Oct 26 09:44:32 2005) Minor failures :
Receiver was trouble.The experiment start record 07:38.we losted scan NO.1 and NO.2. Tsys is high(about 460K).

Metsahovi (Ari Mujunen at Wed Oct 26 10:10:19 2005) Failure : RECE WIND HIGH
Last night we have experienced several massive power outages due to sudden snow, rain, and heavy winds. The K band receiver is warm, it is being cooled again so that it can be "almost-cooled" probably just about in time for EM059A.

Shanghai (An tao at Wed Oct 26 10:24:04 2005) Minor failures : ANT
09:15:39:There has some antenna troubles. 09:36:48.40:we lost the scan no0035 because of antenna problem.

Onsala (Michael Lindqvist at Thu Oct 27 07:07:00 2005) Success :
No known problems.

Torun (Kaz Borkowski at Fri Oct 28 06:21:13 2005) Success :
No known problems.

Torun (Kaz Borkowski at Fri Oct 28 06:22:57 2005) Success :
Sorry - the above was intended for L-band.

Effelsberg (A. Kraus at Fri Oct 28 13:43:49 2005) Success :

Jodrell_Bank (Alastair Gunn at Tue Nov 1 12:17:03 2005) Minor failures : RECO
MkII was used. Communication problems with Mk5 unit at start of experiment. Rebooted and restarted Field System at 08:07 UT. Data good from 08:15 UT. No other known problems.